OpenStudy (anonymous):

A hydrate of CuSO4 has a mass of 11.71 g before heating. After heating, the mass of the anhydrous compound is found to be 8.75 g. Explain how you would determine the formula of the hydrate and then write out the full name of the hydrate.

OpenStudy (anonymous):

Can someone walk me through this? Any help would be greatly appreciated

Parth (parthkohli):

The mass of water of crystallisation = the mass of the initial hydrate - the mass of annhydrate = 11.71 - 8.75 = 1.96 g. The mass of CuSO4 = 8.75 g. Moles of CuSO4 = 8.75/160 Moles of H2O = 1.96/18 You know the quantity of both CuSO4 and H2O. Their simplest ratio would be the formula.